OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y = t^2 + 5

OpenStudy (rajee_sam):

from first eqn you have x = t-3 so t = x + 3 Now put that in the second y = t^2 + 5 put t = x + 3 y = (x + 3)^2 + 5 = x^2 + 6x + 9 +5 = x^2 + 6x + 14

OpenStudy (anonymous):

ohh okay I forgot to square it after I substituted it. thank you could you help me with this one? Determine two pairs of polar coordinates for the point (5, 5) with 0° ≤ θ < 360°

OpenStudy (anonymous):

i got (5 square root 2, 315°), (-5 square root 2, 135°)

OpenStudy (rajee_sam):

I am not good at Polar coordinates

OpenStudy (rajee_sam):

Give me a few minutes I will try

OpenStudy (rajee_sam):

My computer is acting up

OpenStudy (anonymous):

no problem

OpenStudy (rajee_sam):

you have 5 root 2 that is correct. But -5root2 is not. They are asking for all polar coordinates for which magnitude is 5 root 2 and the angle for which tan theta = 1

OpenStudy (rajee_sam):

so your coordinates will be between 0 and 360° \[( 5\sqrt{2}, 45°) ; (5\sqrt{2}, 135° ); (5\sqrt{2}, 225°) ; (5\sqrt{2}, 315°)\]

OpenStudy (anonymous):

which pair is it though?